Circular No. 6801

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ej IN IC 2060 J. Maza, Department of Astronomy, University of Chile, reports the

discovery of a supernova (B ∼ 16.5) by Marina Wischnjewsky on a T-Max 400 film taken by L. González using the Maksutov telescope at Cerro El Roble Observatory on Dec. 29.145 UT. SN 1997ej is located at α = 4^{h}17^{m}53^{s}.49, δ = −56^{o}36′58′′.2 (equinox 2000.0), which is 13′′.5 east and 17′′.5 south of the nucleus of the lenticular galaxy IC 2060 (ESO 157-19). The object is not visible on a IIa-O plate taken with the same telescope on Nov. 25.21, nor is it visible on a T-Max film obtained on Nov. 27.22 (limiting mag ∼ 19.5).

A. Clocchiatti, P. Universidad Catolica de Chile; and M. M. Phillips, Cerro Tololo Interamerican Observatory (CTIO), write: “A spectrogram (range 350–700 nm; resolution 0.6 nm) taken on Dec. 31.1 UT with the CTIO 1.5-m telescope shows that SN 1997ej is a type-Ia supernova probably more than a week before maximum light. The Si II 635.5-nm line is broad and strong, with a flat-bottomed, blueshifted absorption, the centroid of which indicates an expansion velocity of 15 000 km/s. The centroid of the Ca II H and K absorption yields a blueshift of ∼ 19 000 km/s. The spectrum most resembles that of SN 1990N, ∼ 2 weeks before maximum (particularly in the broadness of the spectral features), but it also compares reasonably well with a spectrum of SN 1992bc obtained ∼ 10 days before maximum. The high expansion velocities of the Si II and Ca II lines in SN 1997ej are consistent with such an early phase.”

COMET C/1997 J2 (MEUNIER-DUPOUY)

Ephemeris extension to IAUC 6734 (elements MPC 30738):

1997 /98 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 28 19^{h}12.^{m}86 +41^{o}43.^{′}1 3.404 3.133 65.^{o}9 16.^{o}6 10.6
Jan. 7 19 35.60 +39 44.9 3.443 3.112 62.4 16.3 10.6
17 19 56.85 +37 59.8 3.487 3.094 58.8 15.8 10.6
27 20 16.68 +36 28.7 3.533 3.080 55.2 15.2 10.6
Feb. 6 20 35.13 +35 11.8 3.576 3.068 51.9 14.7 10.6
16 20 52.25 +34 08.4 3.613 3.059 49.1 14.1 10.6
26 21 08.08 +33 17.3 3.641 3.053 47.0 13.7 10.7
Mar. 8 21 22.65 +32 37.4 3.658 3.051 45.9 13.5 10.7
18 21 35.96 +32 06.7 3.660 3.052 45.9 13.5 10.7
28 21 48.02 +31 43.8 3.647 3.056 47.1 13.8 10.7
1997 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green